Trap:

help A die is tossed in an experiment. It lands with the following results: 2, 1, 1, 6, 4, 3, 4, 4, 5, 1. What is the experimental probability of the die landing on an even number?

Graysonunderwoods:

5 out of every 10,time,more or less a 50/50

GRIZZLEY19:

The experimental probability of the die landing on an even number is 0.5 or 50%.

surjithayer:

it is tossed 10 times and even number is 5 times,so experimental probability is 5/10=0.5 or 50%

ihy:

5/10 = 0.5 = 50%

forgetmylife:

@surjithayer wrote:
it is tossed 10 times and even number is 5 times,so experimental probability is 5/10=0.5 or 50%
I agree

toga:

@surjithayer wrote:
it is tossed 10 times and even number is 5 times,so experimental probability is 5/10=0.5 or 50%
I think this is right. @magaback

MAGABACK:

@toga wrote:
@surjithayer wrote:
it is tossed 10 times and even number is 5 times,so experimental probability is 5/10=0.5 or 50%
I think this is right. @magaback
it is indeed. 5/10 is 50%

grimreaper124:

lol